OpenStudy (anonymous):

If the diameter of the base of the cone, d, is 20 units and the height, h, is 19 units, what is the volume of the cone? 300π cubic units 1203.33π cubic units 633.33π cubic units 63.33π cubic units

OpenStudy (anonymous):

Formula for Volume: \[\pi \times r ^{2} \times \frac{ h }{ 3 }\]

OpenStudy (anonymous):

Do you need more help?

OpenStudy (anonymous):

\[\frac{ h }{ 3 }=6.333...\]\[r ^{2}=100\]\[6.333...\times100=?\]That's all I'm doing until I recive an answer from you @Nerd3553

OpenStudy (anonymous):

633.3 ?

OpenStudy (anonymous):

@D.Q.

OpenStudy (anonymous):

Yes, \[633.33\pi\]is correct.

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

Your welcome.